# PISE: Physics-Anchored Semantically-Enhanced Deep Computational Ghost Imaging for Robust Low-Bandwidth Machine Perception

### v1

@@ 1 Introduction
Computational Ghost Imaging (CGI), also known as single-pixel imaging, offers hardware-efficient compressive acquisition [3]. By modulating the light field using structured spatial patterns and measuring aggregated intensity with a bucket detector, the scene can be optically compressed before readout. Recent advances have enabled open-source toolboxes [4] and efficient fiber-laser-based implementations [5] for practical single-pixel imaging systems. However, at deep undersampling (e.g., 5%), the inverse problem becomes massively underdetermined. Traditional reconstruction methods fail in different ways: physics-based solvers exhibit aliasing artifacts [6]; end-to-end deep networks [7, 8, 9, 10] trained with mean squared error (MSE) converge to an overly smoothed conditional mean; and GAN-based methods [11] produce sharp but measurement-inconsistent hallucinations. Deep unfolding approaches [12, 13] have shown promise in balancing interpretability with performance.
Why does MSE training fail under extreme undersampling? This is fundamentally a gradient issue [14]: when many solutions fit the forward model, the MSE gradient decays rapidly. The network converges to a “safe” blurry average that minimizes pixel errors but destroys high-frequency cues required by classifiers. Feature-space loss can restore these cues, but without a physical anchor it may drift towards measurement-inconsistent artifacts.
@@ 2.2.2 Semantic Enhancement via Feature-Space Loss
where $\phi_{j}(\cdot)$ are frozen VGG-16 feature maps [15]. For perceptual loss, grayscale inputs are replicated to 3 channels and ImageNet-normalized before feeding into VGG-16; we use features from layers {relu1_2, relu2_2, relu3_3}. The perceptual term preserves non-zero gradients. We set $\lambda_{\text{mse}}=1.0$ and $\lambda_{\text{perc}}=0.05$. Sensitivity analysis indicates that $\lambda_{\text{perc}}>0.1$ induces oscillatory artifacts, while $\lambda_{\text{perc}}<0.01$ reverts to MSE-like smoothing; 0.05 offers the optimal trade-off.
@@ 3.1 Experimental Setup
Dataset: Fashion-MNIST ($28\times 28$ grayscale apparel images). Sampling: Primary evaluation at $\gamma=5\%$; sensitivity spans $\gamma\in\{2\%,5\%,10\%,20\%\}$. Noise Adaptation Scheme: Training was performed using additive white Gaussian noise (AWGN); evaluation was conducted using Poisson noise to simulate a real detector, and its robustness was tested under mismatched noise conditions. Baselines: ISTA-Net+ [7], ADMM-CSNet [8], U-Net-CS (MSE-only), and cGAN-CS [11]. Metrics: Classification accuracy using the frozen classifier was the primary metric; signal-to-noise ratio (PSNR) was a secondary metric. Loss Weights: We set $\lambda_{\text{mse}}=1.0$. Unless otherwise specified, $\lambda_{\text{perc}}=0.05$ is used in the main comparisons. For sensitivity and ablation studies, $\lambda_{\text{perc}}$ follows the exact values reported in the released configuration files (Supplementary Table S1). We further evaluate on CIFAR-10 natural scenes to validate robustness under photon-limited Poisson noise and 8-bit quantization.While cGAN-CS achieves higher peak accuracy, we prioritize stability and physics consistency for reliable edge deployment.
@@ 3.2 Comparison with Efficient Baselines at 5% Sampling
Figure 2: Visual comparison at $\gamma=5\%$. The adjoint proxy is severely degraded. ISTA-Net+ exhibits aliasing, U-Net (MSE) oversmooths discriminative textures, and cGAN-CS may hallucinate measurement-inconsistent details. PISE recovers semantically faithful structures without spurious artifacts. Note the recovery of high-frequency cues (e.g., shoe laces) compared to blurry MSE results.
@@
cGAN-CS [11] 25.94 16.46 High Variance
@@
Fig. 2 shows the qualitative results. ISTA-Net+ produces highly aliased output. U-Net-CS (MSE) is structurally reasonable but lacks discriminative texture. cGAN-CS is sharp but exhibits semantic hallucinations. PISE recovers readable texture and clear contours without measurement-inconsistent artifacts.
@@ 4 Conclusion
By combining adjoint-based physical anchoring with feature-space semantic guidance, PISE addresses gradient collapse in MSE-based reconstruction under deep undersampling. Minimizing low-order MSE can yield high PSNR but suppress high-frequency cues needed for classifier predictions. Feature-space guidance [15] shifts the objective from pixel means to semantic activation matching, improving recognizability under deep undersampling. Although perceptual loss restores gradients, without constraints it may lead to unstable optimization [14] and measurement-inconsistent artifacts. The adjoint proxy provides a physically grounded anchor that regularizes the search space, improving both accuracy and run-to-run stability.

### v2

@@ 1 Introduction
Computational Ghost Imaging (CGI), also known as single-pixel imaging, offers hardware-efficient compressive acquisition [3]. By modulating the light field using structured spatial patterns and measuring aggregated intensity with a bucket detector, the scene can be optically compressed before readout. Recent advances have enabled open-source toolboxes [4] and efficient fiber-laser-based implementations [5] for practical single-pixel imaging systems. However, at deep undersampling (e.g., 5%), the inverse problem becomes massively underdetermined. Traditional reconstruction methods fail in different ways: classical physics-based / compressive reconstruction methods often yield structured artifacts and degraded detail recovery under low sampling [6]; deep-learning-based ghost imaging methods [7, 8] improve reconstruction quality but remain largely data-driven under severe ambiguity; and adversarial image-translation frameworks [9] can produce visually sharper outputs, motivating the need to better balance semantic fidelity and physical consistency. Deep unfolding approaches [10, 11, 12, 13] have shown promise in balancing interpretability with performance.
Why does MSE training fail under extreme undersampling? In our setting, this manifests as an optimization issue: when many solutions fit the forward model, pixel-wise supervision tends to favor conservative averages and suppress high-frequency cues required by classifiers. Feature-space loss can restore these cues, but without a physical anchor it may drift towards less measurement-faithful structures.
@@ 2.2.2 Semantic Enhancement via Feature-Space Loss
where $\phi_{j}(\cdot)$ are frozen VGG-16 feature maps. For perceptual loss, grayscale inputs are replicated to 3 channels and ImageNet-normalized before feeding into VGG-16; we use features from layers {relu1_2, relu2_2, relu3_3}. The perceptual term preserves non-zero gradients. We set $\lambda_{\text{mse}}=1.0$ and $\lambda_{\text{perc}}=0.05$. Sensitivity analysis indicates that $\lambda_{\text{perc}}>0.1$ induces oscillatory artifacts, while $\lambda_{\text{perc}}<0.01$ reverts to MSE-like smoothing; 0.05 offers the optimal trade-off.
@@ 3.1 Experimental Setup
Dataset: Fashion-MNIST ($28\times 28$ grayscale apparel images). Sampling: Primary evaluation at $\gamma=5\%$; sensitivity spans $\gamma\in\{2\%,5\%,10\%,20\%\}$. Noise Adaptation Scheme: Training was performed using additive white Gaussian noise (AWGN); evaluation was conducted using Poisson noise to simulate a real detector, and its robustness was tested under mismatched noise conditions. Baselines: ISTA-Net+ [10], ADMM-CSNet [11], U-Net-CS (MSE-only), and a pix2pix-style conditional GAN baseline adapted to our CGI setting [9]. Metrics: Classification accuracy using the frozen classifier was the primary metric; signal-to-noise ratio (PSNR) was a secondary metric. Loss Weights: We set $\lambda_{\text{mse}}=1.0$. Unless otherwise specified, $\lambda_{\text{perc}}=0.05$ is used in the main comparisons. For sensitivity and ablation studies, $\lambda_{\text{perc}}$ follows the exact values reported in the released configuration files (Supplementary Table S1). We further evaluate on CIFAR-10 natural scenes to validate robustness under photon-limited Poisson noise and 8-bit quantization.While the pix2pix-style cGAN baseline achieves higher peak accuracy, we prioritize stability and physics consistency for reliable edge deployment.
@@ 3.2 Comparison with Efficient Baselines at 5% Sampling
Figure 2: Visual comparison at $\gamma=5\%$. The adjoint proxy is severely degraded. ISTA-Net+ exhibits aliasing, U-Net (MSE) oversmooths discriminative textures, and the pix2pix-style cGAN baseline may introduce less measurement-faithful details in our setting. PISE recovers semantically faithful structures without obvious spurious artifacts. Note the recovery of high-frequency cues (e.g., shoe laces) compared to blurry MSE results.
@@
pix2pix-style cGAN [9] 25.94 16.46 High Variance
@@
Fig. 2 shows the qualitative results. ISTA-Net+ produces highly aliased output. U-Net-CS (MSE) is structurally reasonable but lacks discriminative texture. The pix2pix-style cGAN baseline is visually sharp but may introduce less measurement-faithful textures in our setting. PISE recovers readable texture and clear contours without measurement-inconsistent artifacts.
@@ 4 Conclusion
By combining adjoint-based physical anchoring with feature-space semantic guidance, PISE addresses gradient collapse in MSE-based reconstruction under deep undersampling. Minimizing low-order MSE can yield high PSNR but suppress high-frequency cues needed for classifier predictions. Feature-space guidance shifts the objective from pixel means to semantic activation matching, improving recognizability under deep undersampling. Although perceptual loss restores gradients, without constraints it may lead to unstable optimization and measurement-inconsistent artifacts. The adjoint proxy provides a physically grounded anchor that regularizes the search space, improving both accuracy and run-to-run stability.